# Collaborative Aquatic Positioning System Utilising Multi-beam Sonar and Depth Sensors

### v2

@@ Abstract
Accurate positioning of remotely operated underwater vehicles (ROVs) in confined environments is crucial for inspection and mapping tasks and is also a prerequisite for autonomous operations. Presently, there are no positioning systems available that are suited for real-world use in confined underwater environments, unconstrained by environmental lighting and water turbidity levels and have sufficient accuracy for long-term, reliable and repeatable navigation. This shortage presents a significant barrier to enhancing the capabilities of ROVs in such scenarios. This paper introduces an innovative positioning system for ROVs operating in confined, cluttered underwater settings, achieved through the collaboration of an omnidirectional surface vehicle and an ROV. A formulation is proposed and evaluated in the simulation against ground truth. The experimental results from the simulation form a proof of principle of the proposed system and also demonstrate its deployability. Unlike many previous approaches, the system does not rely on fixed infrastructure or tracking of features in the environment and can cover large enclosed areas without additional equipment.
@@ I-A Motivation
Over the past few decades, the domain of underwater robotics has experienced considerable expansion. Presently, the use of Remotely Operated Vehicles (ROVs) has become safe and routine, not only within the offshore industry, but also in confined aquatic environments. Numerous examples can be found in the literature, such as pipeline inspection [1], monitoring of nuclear storage facilities [2] and liquid storage tank inspections [3]. Safety concerns and cost-effectiveness are the primary motivations behind this transition from human divers to ROVs [4]. However, navigating underwater vehicles through confined and sometimes cluttered environments is challenging. Precision movement is crucial in these environments, as vehicles must navigate through narrow passages and avoid obstacles [5].
@@
Presently most ROV missions in confined underwater environments are teleoperated [6]. However, there are several benefits to performing fully autonomous missions, including cost reduction, repeatability, and increased survey frequency [7]. Even lower levels of autonomy, such as position and velocity control, can facilitate smooth and accurate navigation or fixing sensor position or velocity for accurate surveying. To achieve autonomy, accurate positioning is essential [6] and currently there are no positioning systems available that are practical for use in confined underwater environments (as discussed in Section II-A). In addition to facilitating autonomy, location-tagged sensor data can be used for long-term environmental monitoring and analysis. In a recent study [8] published from the Fukushima Daiichi Nuclear Power Station in Japan, an investigation was conducted into the environmental conditions of an underground facility used for the temporary storage of contaminated water. Hundreds of sandbags filled with zeolite particles were deployed on the floor to remove radioactivity from the water, with plans for future decommissioning using a zeolite removal robot. The positioning of underwater robots contributed to a more precise survey of the underwater environment and facilitated 3D reconstruction. However, in the current stage of deployment, robots have not addressed the issue of underwater localisation to achieve automation; instead, they are controlled through human intervention.
@@
To be useful in practical situations, a confined space localisation system should require minimal infrastructure, meaning that it should not rely on preinstalled cameras, beacons or sensors. Fixing infrastructure in the environment is time consuming, expensive and generally requires calibration. In addition, installing infrastructure is often not practical. For example, in a nuclear spent fuel pool, where access is highly restricted for safety reasons, it would be impractical to install beacons or other infrastructure around the pool.
@@
In this work, a Collaborative Aquatic Positioning (CAP) system is proposed that aims to satisfy the requirements stated above. Given the prevalent conditions of insufficient illumination and high turbidity in many underwater environments, the proposed localisation system abandons the use of traditional optical cameras for tracking and instead employs multi-beam sonar. The proposed system uses an autonomous surface vehicle (ASV) that can self localise relative to the walls. The surface vehicle tracks and remains above the underwater ROV. By combining sensor measurements from both vehicles, the pose of the subsurface vehicle can be determined. This system requires no fixed infrastructure and minimal calibration. In the next section a review of related underwater localisation systems is provided.
@@ I-B Contribution
    A novel formulation of a collaborative underwater positioning system that is suited for use in real world ROV missions in confined environments and, unlike previous systems with comparable accuracy, does not require any fixed infrastructure, does not rely on tracking environmental, can cover large areas and operate in highly turbid water with no lighting requirements.
@@
    Simulation proof of principle demonstrates the correctness of the proposed CAP-SD mathematical model, showing that the positioning system possesses an acceptable level of accuracy for a variety of applications.
@@ II-A1 ACOUSTIC TRANSPONDERS AND BEACONS
This method includes long baseline (LBL), short baseline (SBL) and ultra short baseline (USBL) systems and is primarily designed for use in open oceans. Baseline localisation is accomplished by measuring distances between transponders and beacons based on the time-of-flight of acoustic signals. Regarding legacy acoustic positioning systems (APS), there are primarily two types: those developed as a result of academic research and those that are commercialised products. In academic research, APS is often integrated with other positioning technologies, such as Doppler Velocity Loggers (DVL) and inertial sensors. In [9], field experiments have demonstrated that for these technologies, the positioning error exceeds 1 meter for the majority of the time. The experimental validation of a tightly coupled USBL and inertial navigation system proposed in [10] also indicates a root mean square estimation error (RMSE) in position ranging between 1 and 2 meters. In commercial APS,  [11] is among the most precise and is typically integrated into medium and small-sized underwater robots, offering positioning accuracy within 0.1 meters for distances up to 5 meters.
@@ II-A2 SIMULTANEOUS LOCALISATION AND MAPPING
Simultaneous localization and mapping (SLAM) is a widely researched topic in the field of robotics for ground and aerial vehicles. However, underwater SLAM is more challenging due to the fact that commonly used sensors such as cameras and LiDARs do not perform well underwater due to higher photon absorption. In [12], Rahman et al. validated three visual SLAM techniques underwater, OKVIS (Stereo) [13], VINS-Mono [14] and ORB-SLAM3 (Stereo-in) [15]. These visual SLAM algorithms can provide localisation with reasonable accuracy: RMSE 0.1-0.3 m. However, visual SLAM needs to constantly identify features in the image to provide accurate positioning, which can be problematic underwater, where visibility is lower than in air and environments are often feature sparse.
@@ II-A4 DEAD RECKONING
Dead reckoning estimates current position based on a previously known or estimated position, using direction and distance traveled. Pure dead reckoning uses exclusively ROV mounted sensors and requires no infrastructure. There are two key technologies that are typically used for underwater dead reckoning: Doppler Velocity Logs (DVLs) [20] and inertial measurement units (IMUs) [21]; these two technologies are commonly combined [22]. However, due to the nature of the method, all dead reckoning systems accumulate errors [23] and are therefore most commonly used to aid other systems that can directly provide position relative to the environment or fixed infrastructure. In [24], an IMU based nonlinear observer for dead reckoning was proposed. It was demonstrated the RMSE of the system was 100 meters after 10 minutes.
@@ II-A5 SUMMARY
It is apparent that although individual technologies meet some of the requirements detailed in Section I-A, none of the technologies reviewed are sufficient to meet the combination of requirements. From a technical perspective, relying solely on multi-beam sonar is insufficient for determining the three-dimensional coordinates of an object [25] [26]. These motivate the development of the CAP system presented in this paper, which integrates a multi-beam sonar and depth sensor (CAP-SD).
@@ III-A Notations and coordinate frames
Figure 1: Overview of the components and coordinate frames of the confined aquatic positioning (CAP-SD) system. An omnidirectional ASV - MallARD, with integrated LiDAR, IMU, and an image sonar tracks a submersible ROV. The ROV is also fitted with a pressure sensor for depth measurement.
Figure 1 presents the coordinate system and reference frames used in this work. The full positioning system is composed of the following co-ordinate frames: the inertial world frame $\mathcal{F}_{W}$, ASV baselink frame $\mathcal{F}_{B}$, multi-beam sonar $\mathcal{F}_{S}$ and ROV frame $\mathcal{F}_{R}$. The world frame $\mathcal{F}_{W}$ origin is assigned to a corner of the testing tank. $\mathcal{F}_{B}$ is the geometric centre of the ASV. The origin of the sonar frame $\mathcal{F}_{S}$ is located at the optical center of the transmitter and has a static transform from $\mathcal{F}_{B}$.
The transformation from one arbitrary coordinate frame $\mathcal{F}_{A}$ to another $\mathcal{F}_{B}$ can be expressed as a homogeneous transformation matrix:
@@
where $\mathbf{R}^{A}_{B}\in\mathrm{SO}(3)$ is a rotation matrix from frame A to frame B and $\mathbf{p}^{A}_{B}\in\mathbb{R}^{3}$ is translation column vector. $\mathbf{H}\in SE(3)$ represents the combination of rotation and translation in three dimensions, which can also be expressed as the pair $[\mathbf{R},\mathbf{p}]$.
III-B CAP-SD formulation overview
The full transform chain from the world frame $\mathcal{F}_{W}$ to the ROV frame $\mathcal{F}_{R}$:
$\mathbf{H}_{S}^{W}=\mathbf{H}_{B}^{W}\mathbf{H}_{S}^{B}\mathbf{H}^{S}_{R},$ (2)
can not be used to obtain $\mathbf{P}^{W}_{R}=[x^{W}_{R},y^{W}_{R},z^{W}_{R}]^{\top}$ directly because $\mathbf{P}^{S}_{R}$ cannot be defined. This is because the multi-beam sonar generates 2-D visualisations of the 3-D environment, where each pixel in the image represents an arc in 3D space, originating from the center of the multi-beam sonar.
In this scenario, each pixel has a direct correlation with both the azimuth angle ($\theta$) and the distance (R) of the arc from its origin, but the aggregation of spatial returns along these arcs precludes the retrieval of elevation angles ($\phi$), as shown in Figure 2A. Therefore, even if the location of the ROV is successfully identified in the sonar’s scan, the translation $\mathbf{P}^{S}_{R}=[x^{S}_{R},y^{S}_{R},z^{S}_{R}]^{\top}$ can not be defined and the transform chain is incomplete. To resolve this problem, the CAP-SD system uses the available elements of the transform chain and the pixel location of the ROV in the sonars image to generate an arc of possible positions of the ROV in the world frame $\mathcal{F}_{W}$. A pressure sensor on the ROV is used to generate a horizontal plane (depth plane) of possible robot locations in $\mathcal{F}_{W}$. By finding the intersection of the arc and the horizontal plane, a single location for the ROV in  $\mathcal{F}_{W}$ can be ascertained.
III-C CAP-SD formulation details
First, the pose of the sonar with respect to world-fixed frame ($\mathcal{F}_{W}$), which also serves as the arc’s center, can be expressed by:
@@ III-C CAP-SD formulation details
where $\mathbf{H}_{S}^{B}$ is a static homogenous transformation from ASV’s baselink frame to sonar frame and $\mathbf{H}_{B}^{W}$ is detailed in Section III-E.
Second, the pixel location of the ROV in the sonar scan is identified using an object detection neural network. The pixel location in the sonar image is used to determine the distance from the sonar to the ROV R and the angle $\theta$, as detailed in Section III-D..
Third, the arc, represented by a circle in 3D space, is defined. This involves first defining a sphere of possible locations and then defining a sphere-cutting plane, where the intersection between the sphere and the sphere-cutting plane defines the circle. The centre of the sphere is the origin of $\mathcal{F}_{S}$ and it’s radius is $R$. The sphere is defined as:
@@
where $[x_{S}^{W}y_{S}^{W}z_{S}^{W}]^{\top}$ is the position of multi-beam sonar. The sphere-cutting plane can be defined by taking the plane of the $x-z$ axes of $\mathcal{F}_{S}$ and rotating it by the angle $\theta$ about the $y$ axis of $\mathcal{F}_{S}$ (gray plane depicted in Figure 2B). The sphere-cutting plane containing the arc can be represented as:
@@
where $\vec{\mathbf{n}}_{3\times 1}=[A,B,C]^{\top}$ is the normal vector to the cutting-sphere plane, defined as:
$\left[\begin{array}[]{c}\vec{\mathbf{n}}_{3\times 1}\\ 1\end{array}\right]=(H^{W}_{B}H^{B}_{S})^{\prime}\left[\begin{array}[]{c}\sin(\theta)\\ -\cos(\theta)\\ 0\\ 1\end{array}\right].$ (6)
Fourth, the depth plane in $\mathcal{F}_{W}$ must be defined. This is achieved by attaching a pressure sensor to the ROV and calibrating it to measure depth below the surface. The depth plane, which is normal to the $\mathcal{F}_{W}$ $z$-axis, can be defined as:
@@
where $z_{depth}$ is the depth measured by the pressure sensor.
Fifth, by intersecting the sphere, sphere-cutting plane, and the depth plane, the solution domain can be reduced to two points in $\mathcal{F}_{W}$. This is achieve by solving equations (4), (5) and (7) for $x,y$ and $z$.
The final step is to ascertain which of the two points is the correct one. This is relatively simple as the correct solution lies in the positive $x$ region in $\mathcal{F}_{S}$, as shown orange arc in Figure 2C). When transformed into the world-fixed frame, the $x^{W}_{R}$ and $y^{W}_{R}$ must satisfy the following condition:
$R_{11}x^{W}_{R}+R_{12}y^{W}_{R}+R_{13}z^{W}_{R}+t_{x}>0,$ (8)
where $R_{11}$, $R_{12}$, $R_{13}$ and $t_{x}$ are elements of the transform:
$H^{W}_{S}=\left(\begin{array}[]{cccc}R_{11}&R_{12}&R_{13}&t_{x}\\ R_{21}&R_{22}&R_{23}&t_{y}\\ R_{31}&R_{32}&R_{33}&t_{z}\\ 0&0&0&1\\ \end{array}\right).$ (9)
Therefore, a single point is ultimately identified, which represents the robot’s position $\mathbf{P}^{W}_{R}=[x^{W}_{R},y^{W}_{R},z^{W}_{R}]^{\top}$ with respect of the world-fixed coordinate frame $\mathcal{F}_{W}$.
Figure 2: Abstract frames of the CAP-SD system. (A) Underwater side-view of CAP-SD system. (B) Using the cutting-sphere plane and sphere to define the arc where the ROV may be located. (C) Possible locations of ROV can be reduced to two points ($P_{1}$ and $P_{2}$) by generating another plane using the depth sensor. The position of ROV ($P_{2}$) must be in the sonars positive x region (orange arc in (B) and (C)).
III-D Neural network (NN) for image sonar object detection
Figure 3: NN object detector (A) Side view of a working scenario involving multi-beam sonar (B)  Raw sonar image (C) Sonar image within NN object detector (D) Pixel array and sonar image coordinate interchange
The NN detector serves to pinpoint the robot’s pixel coordinates within the sonar image. Given the predefined horizontal field-of-view (FOV), maximum range, and image resolution of an actual multi-beam sonar, determining the robot’s pixel coordinates in the sonar image consequently specifies its distance $R$ and azimuth angle $\theta$ in relation to the multi-beam sonar.
A pre-trained YOLOv5 model  [^1] The detail about the customized pre-trained model can be found in: https://github.com/Xue1iang/CAP_SD_Unity was employed to detect robots in sonar images, which is does by identifying them within bounding boxes as shown in Fig 3C. YOLOv5 builds on the previous YOLO version, with improvements introduced to further boost performance and flexibility to ease deployability in real-world experiments. The pixel coordinates of the robot are obtained at this stage. Subsequently, a conversion from pixel coordinates to sonar image coordinates is performed, as depicted in Fig 3D. For the sonar image frame, its coordinate parameters, such as range and vertical width, are predetermined and fixed. Therefore, pixel coordinates can be directly converted into the sonar image’s $\left\{R,\theta\right\}$ coordinates.
III-E Calculating ASV rotation relative to the world frame $\mathbf{R}^{W}_{B}$
To localise the surface robot within the world-fixed frame, a homogeneous transformation, $\mathbf{H}^{W}_{B}=[\mathbf{R}^{W}_{B},\mathbf{p}^{W}_{B}]$, is calculated. Utilizing a tilting extended Kalman filter (EKF), IMU and SLAM data are fused to estimate the Unmanned Surface Vehicle’s (USV) orientation, $\mathbf{R}^{W}_{B}$. This process incorporates tilt angles from the IMU and the yaw angle from SLAM, adopting a Z-Y-X Euler sequence for the comprehensive rotation matrix construction. Despite Euler angles’ susceptibility to gimbal lock, such occurrences are mitigated here by ensuring the $y$-axis rotation does not exceed 90 degrees, a condition met by surface vehicles.
The EKF is used to accurately estimate pitch and roll angles ($\psi^{W}_{B},\theta^{W}_{B}$) under dynamic conditions, utilizing data from a three-axis gyroscope and accelerometer. Inputs to the tilting EKF include tri-axial angular rates, $\boldsymbol{\omega}=[{\omega}_{x},{\omega}_{y},{\omega}_{z}]^{\top}$, and accelerations, $\boldsymbol{a}=[{a}_{x},{a}_{y},{a}_{z}]^{\top}$, from the USV’s gyroscope and accelerometer, respectively, with the assumption of uncorrelated Gaussian noise. The model assumes negligible non-gravitational linear accelerations during the USV’s motion.
The position and orientation data $[x^{W}_{B},y^{W}_{B},\psi^{W}_{B}]^{\top}$ provided by SLAM (Hector SLAM [27] was deployed on the ASV in the simulation) is notably accurate, as verified by our prior research [28]. The fusion of SLAM and IMU data results in Euler angles $[\psi^{W}_{B},\theta^{W}_{B},\phi^{W}_{B}]^{\top}$, with pitch ($\theta$) and roll ($\phi$) derived from the tilting EKF. Consequently, the rotation matrix $\mathbf{R}^{W}_{B}$ is formulated using the conventional Z-Y-X Euler angles based on the fused orientation data.

### v3

@@ Abstract
Accurate positioning of underwater robots in confined environments is crucial for inspection and mapping tasks and is also a prerequisite for autonomous operations. Presently, there are no positioning systems available that are suited for real-world use in confined underwater environments, unconstrained by environmental lighting and water turbidity levels, and have sufficient accuracy for reliable and repeatable navigation. This shortage presents a significant barrier to enhancing the capabilities of remotely operated vehicles (ROVs) in such scenarios. This paper introduces an innovative positioning system for ROVs operating in confined, cluttered underwater settings, achieved through the collaboration of an omnidirectional surface vehicle and an underwater ROV. A mathematical formulation based on the available sensors is proposed and evaluated. Experimental results from both a high-fidelity simulation environment and a mock-up of an industrial tank provide a proof of principle for the system and demonstrate its practical deployability in real-world scenarios. Unlike many previous approaches, the system does not rely on fixed infrastructure or tracking of features in the environment and can cover large enclosed areas without additional equipment.
@@ I-A Motivation
Over the last decade, the domain of underwater robotics has experienced considerable expansion. Presently, the use of ROVs has become safe and routine, not only within the offshore industry, but also in confined aquatic environments. Numerous examples of their successful deployment can be found in the literature, such as pipeline inspection [1], monitoring of nuclear storage facilities [2] and liquid storage tank inspections [3]. Safety concerns and cost-effectiveness are the primary motivations behind the transition from human divers to ROVs [4]. However, navigating underwater vehicles through confined and sometimes cluttered environments is challenging. Precision movement is crucial in these environments, as vehicles must often navigate through narrow passages and avoid obstacles [5].
@@
Presently most ROV missions in confined underwater environments are teleoperated [6]. However, there are several benefits to performing fully autonomous missions, including cost reduction, repeatability, and increased survey frequency [7]. Aside from full autonomy, lower levels of autonomy, such as position and velocity control, can facilitate smooth and accurate navigation or accurate logging of sensor positions and trajectories for accurate surveying. To achieve autonomy at any level, accurate positioning is essential [6] and currently there are no positioning systems available that are practical for use in confined underwater environments (as discussed in Section II-A). In addition to facilitating autonomy, location-tagged sensor data can be used for long-term environmental monitoring and analysis. In a recent study [8] published from the Fukushima Daiichi Nuclear Power Station in Japan, an investigation was conducted into the environmental conditions of an underground facility used for the temporary storage of contaminated water. Hundreds of sandbags filled with zeolite particles were deployed on the floor to remove radioactivity from the water, with plans for future decommissioning using a zeolite removal ROV. The positioning of ROVs contributed to a more precise survey of the underwater environment and facilitated 3D reconstruction. However, in the current stage of deployment, ROVs have not addressed the issue of underwater localisation to achieve automation and instead, they are controlled through human intervention.
@@
To be useful in practical situations, a confined space localisation system should require minimal infrastructure, meaning that it should not rely on preinstalled cameras, beacons or sensors. Fixing infrastructure in the environment is time consuming, expensive and generally requires calibration. In addition, installing infrastructure is often not practical. For example, in a nuclear spent fuel pool, where access is highly restricted for safety reasons, it is impractical to install beacons or other infrastructure around the pool.
@@
In this work, a Collaborative Aquatic Positioning (CAP) system is proposed that aims to satisfy the requirements stated above. Given the prevalent conditions of insufficient illumination and high turbidity in many underwater environments, the proposed localisation system replaces the use of optical cameras for tracking and instead employs multi-beam sonar. The proposed system uses an autonomous surface vehicle (ASV) that can self localise relative to the walls. The surface vehicle tracks and remains above the underwater ROV. By combining sensor measurements from both vehicles, the pose of the subsurface vehicle can be determined. The resulting system requires no fixed infrastructure and minimal calibration. In the next section a review of related underwater localisation systems is provided.
@@ I-B Contribution
    A novel formulation of a collaborative underwater positioning system that is suited for use in real world ROV missions in confined environments is presented. Unlike previous systems with comparable accuracy, the proposed system does not require any fixed infrastructure, does not rely on tracking environmental features, can cover large areas, and can operate in highly turbid water with no lighting requirements.
@@
    Both simulation and proof of principle studies demonstrate the correctness and effectiveness of the proposed CAP-SD mathematical model, confirming that the positioning system possesses an acceptable level of accuracy for a variety of applications.
@@ II-A1 ACOUSTIC TRANSPONDERS AND BEACONS
This method includes long baseline (LBL), short baseline (SBL) and ultra short baseline (USBL) systems and is primarily designed for use in open oceans. Baseline localisation is accomplished by measuring distances between transponders and beacons based on the time-of-flight of acoustic signals. Regarding legacy acoustic positioning systems (APS), there are primarily two types: those developed as a result of academic research and those that are commercialised products. In academic research, APS is often integrated with other positioning technologies, such as Doppler Velocity Loggers (DVL) and inertial sensors. In [9], field experiments have demonstrated that for these technologies, the positioning error exceeds 1 meter for the majority of the time. The experimental validation of a tightly coupled USBL and inertial navigation system proposed in [10] also indicates a root mean square estimation error (RMSE) in position ranging between 1 and 2 meters. In commercial APS,  [11] is among the most precise and is typically integrated into medium and small-sized underwater ROVs, offering positioning accuracy within 0.1 meters for distances up to 5 meters.
@@ II-A2 SIMULTANEOUS LOCALISATION AND MAPPING
Simultaneous localization and mapping (SLAM) is a widely researched topic in the field of robotics for ground and aerial vehicles. However, underwater SLAM is more challenging due to the fact that commonly used sensors such as cameras and LiDARs do not perform well underwater due to higher photon absorption. In [12], Rahman et al. validated three visual SLAM techniques underwater, OKVIS (Stereo) [13], VINS-Mono [14] and ORB-SLAM3 (Stereo-in) [15]. Their work demonstrated that the visual SLAM algorithms could provide localisation with reasonable accuracy: RMSE 0.1-0.3 m. However, visual SLAM needed to constantly identify features in the image to provide accurate positioning, which is problematic in underwater environments, where visibility is lower than in air and environments are often feature sparse.
@@ II-A4 DEAD RECKONING
Dead reckoning estimates current position on a previously known or estimated position, and then using the current direction and distance travelled to update this position. Pure dead reckoning uses exclusively ROV mounted sensors and requires no infrastructure. There are two key technologies that are typically used for underwater dead reckoning: Doppler Velocity Logs (DVLs) [20] and inertial measurement units (IMUs) [21]; these two technologies are commonly combined [22]. However, due to the nature of the method, all dead reckoning systems accumulate errors [23] and are therefore most commonly used to aid other systems that can directly provide position relative to the environment or fixed infrastructure. In [24], an IMU based nonlinear observer for dead reckoning was proposed. It was demonstrated the RMSE of the system was 100 meters after 10 minutes.
@@ II-A5 SUMMARY
It is apparent that although individual technologies meet some of the requirements detailed in Section I-A, none of the technologies reviewed are sufficient to meet the combination of requirements. From a technical perspective, relying solely on a multi-beam sonar is insufficient for determining the three-dimensional coordinates of an object [25] [26]. These motivate the development of the CAP system presented in this paper, which integrates a multi-beam sonar and depth sensor (CAP-SD).
@@ III-A Notations and coordinate frames
Figure 1: Overview of the proposed CAP-SD system. An omnidirectional ASV - MallARD, with integrated LiDAR, IMU, and an image sonar tracks a submersible ROV. The ROV is also fitted with a pressure sensor for depth measurement.
Figure 1 presents the coordinate system and reference frames used in this work. The complete positioning system is composed of the following co-ordinate frames: world frame, $\mathcal{F}_{W}$, ASV baselink frame, $\mathcal{F}_{B}$, multi-beam sonar frame, $\mathcal{F}_{S}$ and underwater ROV frame, $\mathcal{F}_{R}$. The origin of  $\mathcal{F}_{W}$ has been assigned to a corner of the testing tank. $\mathcal{F}_{B}$ is the geometric centre of the ASV and  $\mathcal{F}_{S}$ has a fixed transform from $\mathcal{F}_{B}$.
The transformation from one arbitrary coordinate frame, $\mathcal{F}_{A}$, to another, $\mathcal{F}_{B}$, can be expressed as a homogeneous transformation matrix:
@@
where $\mathbf{R}^{A}_{B}\in\mathrm{SO}(3)$ is a rotation matrix from frame A to frame B and $\mathbf{p}^{A}_{B}\in\mathbb{R}^{3}$ is a translation column vector. $\mathbf{H}\in SE(3)$ represents the combination of rotation and translation in three dimensions, which can also be expressed as the pair $\mathbf{R},\mathbf{p}]$.
III-B Estimating the sonar pose in the world frame
The proposed CAP-SD system initially performs self-localisation of the ASV by combining 2D SLAM with IMU (Inertial Measurement Unit) data, to ascertain the pose of the ASV relative to the world-fixed frame. Details of the method for doing this have been presented in a previous paper [27] by the author, with an overview provided below.
The position and orientation data $x^{W}_{B},y^{W}_{B},\psi^{W}_{B}]^{\top}$ provided by SLAM is notably accurate, as verified by our prior research [28]. The fusion of SLAM and IMU data results in Euler angles $\psi^{W}_{B},\theta^{W}_{B},\phi^{W}_{B}]^{\top}$, with pitch ($\theta$) and roll ($\phi$) derived from a tilting EKF. Consequently, the rotation matrix, $\mathbf{R}^{W}_{B}$, is formulated using the conventional Z-Y-X Euler angles based on the fused orientation data.
To localise the ASV within the world-fixed frame, a homogeneous transformation, $\mathbf{H}^{W}_{B}=[\mathbf{R}^{W}_{B},\mathbf{p}^{W}_{B}]$, is calculated. Utilising a tilting EKF, IMU and SLAM data are fused to estimate the USV’s orientation, $\mathbf{R}^{W}_{B}$. This process incorporates tilt angles from the IMU and the yaw angle from SLAM, adopting a Z-Y-X Euler sequence for the comprehensive construction of the rotation matrix. Despite the susceptibility of Euler angles to gimbal lock, such occurrences are mitigated here by ensuring the $y$-axis rotation does not exceed 90 degrees, a condition met by surface vehicles.
The EKF was used to accurately estimate pitch and roll angles ($\psi^{W}_{B},\theta^{W}_{B}$) under dynamic conditions, utilising data from a three-axis gyroscope and accelerometer. Inputs to the tilting EKF include tri-axial angular rates, $\bm{\omega}=[{\omega}_{x},{\omega}_{y},{\omega}_{z}]^{\top}$, and accelerations, $\bm{a}=[{a}_{x},{a}_{y},{a}_{z}]^{\top}$, from the USV’s gyroscope and accelerometer, respectively, with the assumption of uncorrelated Gaussian noise. The model assumes negligible non-gravitational linear accelerations during the USV’s motion.
The pose of the sonar with respect to ($\mathcal{F}_{W}$) can be expressed by:
@@ III-B Estimating the sonar pose in the world frame
III-C Identifying underwater ROV’s pixels in sonar scans
Figure 2: NN object detector (A) Side view of a working scenario involving the commercially available BlueROV ROV, equipped with multi-beam sonar (B)  Raw sonar image (C) Sonar image within NN object detector (D) Pixel array and sonar image coordinate interchange
A NN detector serves to pinpoint the ROV’s pixel coordinates within the sonar image. Given the predefined horizontal field-of-view (FOV), maximum range, and image resolution of an actual multi-beam sonar, determining the ROV’s pixel coordinates in the sonar image consequently specifies its distance $R$ and azimuth angle $\theta$ in relation to the multi-beam sonar.
A pre-trained YOLOv5 model  [^1] The detail about the customized pre-trained model can be found in: https://github.com/Xue1iang/CAP_SD_Unity was employed to detect the locations of any ROVs in the sonar images. This is achieved by identifying any ROVs within bounding boxes as shown in Fig 2C. YOLOv5 builds on the previous YOLO version, with improvements introduced to further boost performance and flexibility to ease deployability in real-world experiments. The pixel coordinates of any ROV are obtained at this stage. Subsequently, a conversion from pixel coordinates to sonar image coordinates is performed, as depicted in Fig 2D. For the sonar image frame, its coordinate parameters, such as range and vertical width, are predetermined and fixed. Therefore, pixel coordinates can be directly converted into the sonar image’s $\left\{R,\theta\right\}$ coordinates.
III-D Estimate underwater ROV position in world frame
Figure 3: Abstract frames of the CAP-SD system. (A) Underwater side-view of CAP-SD system. (B) Using the cutting-sphere plane and sphere to define the circle where the ROV may be located. (C) Possible locations of ROV can be reduced to two points ($P_{1}$ and $P_{2}$) by generating another plane using the depth sensor. The position of ROV ($P_{2}$) must be in the sonar’s positive x region (orange arc in (B) and (C)).
Based on $H^{W}_{S}$ mentioned in Section III-B and $\left\{R,\theta\right\}$ discussed in Section III-C, a circle of possible positions for the ROV can be determined by solving the intersection between a sphere and a cutting-sphere plane:
@@ III-D Estimate underwater ROV position in world frame
where $x_{S}^{W}$, $y_{S}^{W}$ and $z_{S}^{W}$ is the position of multi-beam sonar. The cutting-sphere plane containing the circle can be represented as (gray plane depicted in Figure 3B):
@@
where $\vec{n}_{3\times 1}=[A,B,C]^{\top}$ is the normal vector to the cutting-sphere plane, defined as:
$\left[\begin{array}[]{c}\vec{n}_{3\times 1}\\ 1\end{array}\right]=(H^{W}_{B}H^{B}_{S})^{\prime}\left[\begin{array}[]{c}\sin(\theta)\\ -\cos(\theta)\\ 0\\ 1\end{array}\right].$ (5)
To determine the position of the ROV in  $\mathcal{F}_{W}$, it is necessary to utilise a depth sensor that can define a $z$ plane in $\mathcal{F}_{W}$ :
@@
The plane (purple plane in Figure 3C) defined by the depth sensor, cuts through the circle (gray circle in Figure 3C) of possible ROV positions, resulting in two intersection points. However, the ROV that is to be located, must appear within the FOV of the image sonar (sonar’s positive x region as shown in the orange arc in Figure 3C). Therefore, a single point is ultimately identified, which represents the position of the ROV $\mathbf{P}^{W}_{R}=[x^{W}_{R},y^{W}_{R},z^{W}_{R}]^{\top}$ with respect to  $\mathcal{F}_{W}$.